Show that the escape velocity from the surface of earth is √2 times the velocity of projection of an artificial satellite orbiting close around the earth.


1
Expert's answer
2021-03-04T11:57:01-0500

The escape velocity from the surface of earth is


"v_e=\sqrt{\frac{2GM}{R}}"

For orbiting around the earth:


"v=\sqrt{\frac{GM}{R+h}}"

For a satellite very close to Earth


"v=\sqrt{\frac{GM}{R+0}}=\sqrt{\frac{GM}{R}}"

Thus


"\frac{v_e}{v}=\sqrt{2}"